OpenStudy (anonymous):

Use the quadratic formula to solve 6x2 - 17x + 12 = 0

OpenStudy (mrnood):

\[6x ^{2}-17x+12=0\] \[ax ^{2}+bx+c=0\] You know the quadratic formula. so from the above illustration what are a, b and c? Substitute a, b and c into the formula and you will get the two values for x The quadratic equation is there to help YOU solve quadratic equations....

OpenStudy (mrnood):

the formula is HERE: http://0.tqn.com/d/create/1/0/i/9/9/-/Quadratic.formula.jpg

OpenStudy (anonymous):

I don't know the quadratic equation.

OpenStudy (mrnood):

The link above gives you the equation

OpenStudy (anonymous):

the answer I got was x=3/2 or 4/3 is that correct?

OpenStudy (mrnood):

I don't know - the help above allows you to work it out. If you want to check your answer the put the values back into the equation.. \[6(\frac{ 3 }{ 3 })^{2}-17(\frac{ 3 }{ 2 })+12\] If that is equal to 0 then one of your answers is correct. You can check the other one the same way

OpenStudy (mrnood):

(should be 3/2 above - not 3/3)